OpenStudy (tiffany_rhodes):

Can you help me please? Find the area of the region between the curves y=abs(x) and y=(x^2)-2

OpenStudy (anonymous):

Start with the limits of integration for the integral. For the limits of integration, we find where \[\left| x \right|=x^{2}-2\]

jimthompson5910 (jim_thompson5910):

I would take advantage of the symmetry with the y axis. Both equations are symmetric with respect to the y axis.

OpenStudy (anonymous):

They intersect at x=-2 and x=2, so we do:\[\int\limits_{-2}^{2}(\left| x \right|-(x ^{2}-2))dx\] abs(x) is first because abs(x)>x^2-2

OpenStudy (anonymous):

on that interval

OpenStudy (anonymous):

And yeah we could do it from 0 to 2 and multiply that by 2

OpenStudy (tiffany_rhodes):

okay, so I integrated from 0 to 2 and then multiplied by 2 to find the area on the other side. My final answer was 20/3 and it was correct. Thank you both so much!

OpenStudy (anonymous):

Sure